tickets to a concert are 5 dollars for balcony and 10 for orchestra. if the attendance at one show was 800 and if receipts were $700, how many people bought orchestra seats

Answer by Boreal(15235)

Right from the beginning, this can't be solved.
Attendance is 800, and even if all the tickets were for the balcony only, the receipts would be $4000. I will edit to $7000 receipts, instead of $700.
.
concert=x
orchestra= 800-x
5x+10(800-x)=7000
5x+8000-10x=7000
-5x=-1000
x=200
800-x=600
Check
200*5=1000
600*10=6000
This works